OpenStudy (anonymous):

Find the slope of a line that is perpendicular to the line containing the points (-2, -1) and (2, -3) m= -2 m= 2 m= -1/2 m= -1

OpenStudy (texaschic101):

lets first find the slope slope = (y2 - y1) / (x2 - x1) (-2,-1) x1 = -2 and y1 = -1 (2,-3) x2 = 2 and y2 = -3 now we sub slope = (-3 - (-1) / (2 - (-2) slope = (-3 + 1) / (2 + 2) slope = -2/4 slope = -1/2 we are looking for a perpendicular line, so we need the negative reciprocal slope. All that means is " flip " the slope and change the sign. So the negative reciprocal of -1/2 is 2/1, or just 2 (see how I flipped the slope and changed the sign) The slope for the perpendicular line is 2

OpenStudy (mathmale):

Hint: You can now pick either of the given two points. Identify its coordinates (x- and y-values). As texaschic101 has already demonstrated, m = 2. Substitute those x- and y-coordinates from your ONE point, along with your m = 2, into the "point-slope" formula for the eq'n of a straight line:\[y-y _{1}=m(x-x _{1})\] and then simplify your result.

OpenStudy (anonymous):

thank you guys so much!(: